On June 15, 2021 Mars Company owns 10,000 shares with a cost P 700,000 of Moon Company’s stocks. During the same period, Moon Company right to existing shareholder. Mars received 10,000 stock rights entitling him to purchase 5,000 new shares at P 80. The ordinary share was trading ex rights. On July 15,2021 mars exercised all the stocks rights. The share is quoted right on at P90.

1. Assuming that the above securities are FVTPL, the cost of the investment acquired through exercised of stocks rights should be

a. Nil c. P600,000

b. P400,000 d. None of the choices

2. Assuming that the above securities are FVTOCI, the cost of the investment acquired through exercised of stocks rights should be

a. Nil c. P600,000

b. P400,000 d. None of the choices



1
Expert's answer
2021-06-07T11:30:01-0400

1.d

"5000\times 80+5000(90-80)=450 000"


2.d

"5000\times 80+5000\times90=850 000"